### v1

@@ 3.1 Medical Adversarial Defensive Strategies
Various defense models, including techniques such as input denoising, input gradients regularization, and adversarial training, have been devised. However, recent attacks often manage to evade or partially circumvent these protective measures. This paper undertakes a comprehensive exploration of adversarial attack and defense methods within the domain of medical image analysis. It introduces a novel taxonomy based on the application context for a family of techniques, as outlined in Dong et al.’s work [45]. The author establishes a unified theoretical framework encompassing diverse adversarial attack and defense strategies tailored for medical imaging applications.
The primary emphasis in most research efforts on adversarial attacks in medical image processing is on the white-box scenario. These studies, characterized by a comprehensive understanding of medical Deep Neural Networks (DNNs), concentrate on the susceptibility of computer-aided diagnosis models across various medical imaging applications. In the execution of adversarial attacks, attackers may utilize the target diagnosis DNN as a locally deployed model. In addition to white-box adversarial attacks for medical classification tasks, scholars have delved into vulnerabilities associated with various medical imaging tasks. Notably, these publications focus predominantly on medical segmentation, as evident in works such as Chen et al.’s [35] and Ozbulak et al.’s [32].
For natural photos, semi-white-box (Gray-box) attacks have been extensively researched [46]. However, there are only a few academic publications [47], [48] that address this attack scenario for medical image processing. The semi-white-box adversarial approach typically consists of two stages: 1) To create adversarial instances against the target DNN model, the attacker trains a generative model. The attacker has complete access to the target model during training, including backward propagation gradients. 2) Instead of needing to know anything about the target model, as would be the case in a completely black-box scenario, the adversary generator can directly obtain adversarial examples against the target model with the input of authentic photos during the application stage.
Presently utilized white-box adversarial attacks often require multiple backward gradients of the target model. In simpler terms, attackers generate comparable adversarial samples by treating the target Deep Neural Network (DNN) as if it were a locally deployed model. However, due to its reliance on an in-depth understanding of the DNN model to execute an attack, the white-box approach may be unreliable in real-world scenarios. Conversely, the general black-box scenario may offer a more suitable environment for simulating real-world adversarial attacks, with numerous proposals focusing on investigating black-box assaults for natural images.
The outlined solutions usually involve multiple queries to the black-box model or depend on having a comprehensive understanding of the desired diagnosis model. However, in many real-world situations, attackers may lack direct access to the target diagnosis model. The restricted black-box (no-box) configuration, specifically representing the most challenging scenario for real-world adversarial assaults, may be more covertly dangerous even without querying the target black-box DNN. The transferability [49] of adversarial instances across different DNN models fundamentally determines the success of no-box attacks. For example, a no-box attacker can generate adversarial images based on a locally deployed surrogate model, transferring them directly to target medical diagnosis systems. Restricted black-box adversarial attacks are deemed more subtly perilous for natural vision tasks according to a renowned study.
Considering the substantial risks to the healthcare sector, numerous defense strategies have been proposed to counter medical adversarial attacks. In order to overcome the weaknesses shown by adversarial examples, a number of methods have been developed to provide trustworthy deep learning-based systems for natural pictures [49]. The development of accurate computer-aided diagnosis models for clinical applications not only contributes to trustworthy healthcare services for millions but also underscores the importance of investigating adversarially robust models within the realm of medical image analysis.
In their work [45], the authors succinctly summarize studies on adversarial defense in the context of medical image analysis. Various attack methods in the realm of medical images serve as robustness evaluation criteria for adversarial defense, complementing the extensive range of adversarial attack methods mentioned. The minimal separation between adversarial examples and the model’s decision boundary signifies the model’s resistance to disturbances. The goal of adversarial training is to mitigate subsequent losses.
Minimize the
$L(f(x_{\text{adv}},y)+\lambda\cdot L(x,x_{\text{adv}}))$ (1)
where $L(f(x_{\text{adv}},y)$ is the classification loss, $L(x,x_{\text{adv}})$ is the perturbation size, and $\lambda$ is a trade-off hyperparameter.
Transforming the input data into an alternative feature space that diminishes the impact of adversarial perturbations on the model. A technique aimed at enhancing resilience against adversarial attacks involves introducing noise to the features or intermediate representations that characterize the input data.
@@ 3.3 Adversarial Detection
Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models.Many adversarial detection techniques have been put forth in the field of medical image analysis to stop additional misdiagnosis caused by adversarial samples [42]. In particular, it is possible to think about medical adversarial detection as an anomaly detection issue that can be resolved by combining explainability approaches [53].
@@ 3.4 Image-level Pre-processing
3.5 Improvement of Features
The difference in robustness between human and machine vision is attributed to adversarial examples linked to non-robust features extracted from specific patterns in the data distribution [54]. Consequently, enhancing feature representation is crucial for the development of robust inference systems. In this investigation, we characterize feature augmentation as the modification of architectures or mapping functions. Various techniques aimed at improving features have significantly enhanced the robustness of medical classification models [55].
@@ 3.6 Distillation of Knowledge
In the field of machine learning, knowledge distillation is a helpful technique for transferring learned information from a complex (teacher) model to a simple (student) model. Therefore, the situation where the network structures for the teacher and learners are identical is specifically referred to as self-distillation. Furthermore, a great deal of research has been done on the natural imaging domain using adversarial knowledge distillation, which moves the adversarial resilience from a heavy teacher model to a view student model [56].
@@ 4 Result and Discussion
This study utilizes four publicly accessible benchmark datasets, as detailed in [45], to investigate adversarial attack and defense in the context of medical image processing (illustrated in Figure 6). Firstly, the Messidor1 dataset comprises 1,200 eye fundus color images for detecting diabetic retinopathy across four classes based on retinopathy grade. Secondly, the International Skin Imaging Collaboration (ISIC) dataset consists of 2,750 dermoscopic images categorized into three classes for skin lesion classification and segmentation. Thirdly, the ChestX-ray 14 dataset comprises 112,120 frontal-view X-ray images representing 14 thorax diseases. Lastly, the COVID-19 database incorporates 21,165 chest X-ray images accompanied by segmentation-capable lung masks.
@@
The study conducted by the authors [57] employs experiments to demonstrate the following: 1) The SSAT module significantly enhances the adversarial robustness of the model without compromising the classification accuracy of clean images. 2) The UAD module effectively identifies and excludes a majority of successful adversarial examples. 3) In comparison to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes the risk of adversarial attacks. The research utilizes a publicly available dataset of retinal OCT images, employing the ”white-box” setting as the most challenging threat for assessing class prediction performance. Across all white-box attack conditions, the authors establish that SSAT consistently outperforms alternative baselines while maintaining comparable or superior performance for clean image classification. Furthermore, the authors demonstrate that the combined approach of UAD complementing SSAT results in the lowest adversarial risk according to the new measure presented. Regardless of the training techniques employed, systems based on UAD consistently exhibit reduced risks compared to those that do not incorporate this defense mechanism.
@@
An AI-CAD model depicted in Figure 7 was initially trained to manipulate diagnosis-sensitive elements within images, such as adding or removing malignant tissue. The model’s effectiveness was then assessed using adversarial images produced by a GAN model. To determine how well human specialists could visually distinguish images produced by the GAN, a reader study was carried out.
In a study by Zhou et al. [58], the responses of an AI-CAD model to adversarial attacks on GAN-generated mammography images were investigated. This involved introducing malignant tissue into healthy images and removing cancerous tissue from images affected by cancer. The study also evaluated the proficiency of experienced radiologists in visually identifying such adversarial images, both before and after instructional intervention. The study cohort, provided by the University of Pittsburgh Medical Center, included 1284 women, and 4346 mammography images were collected. Among these, 366 patients had biopsy-proven malignant breast cancer, while 918 patients had breast cancer assessed as benign (including benign signs).
Following training, the researchers assessed the model’s classification accuracy concerning both the original genuine test data and its equivalent GAN-generated adversarial counterparts. Two GAN-trained U-Net23 models were employed to generate adversarial images, flipping labels in the test set. The high-resolution (1728 x 1408) photographs indicated the categorization impact of the AI-CAD model on the test data. The AI-CAD model achieved an AUC of 0.82 on the test set, consisting of 364 genuine negative samples and 74 real positive samples. The adversarial GAN-generated images (with flipped labels) in the test set yielded an AUC of 0.94. These AUC values suggested that the adversarial collection of images effectively deceived the AI-CAD model. Additionally, 59.5% (44 out of 74 cases) of genuine positive images with a classification accuracy threshold of 0.5 were correctly labeled as positive, whereas 95.5% (42 out of 44 cases) of adversarial samples created by GANs successfully tricked the system by being classified as negative.
@@
Figure 8: The Mean Structural Similarity Index Measure (SSIM) was computed between the authentic images in the test sets and the adversarial instances generated using FGSM or PGD at different perturbation levels. The SSIM values displayed represent the average across two model architectures (Inception-v3 and Densenet-121) [59].
@@
Figure 8 displays the mean SSIM values for FGSM and PGD assaults across all photos. The Supplementary Material includes the SSIM data for every model. As can be shown, the effects of the identical disturbance applied to various imaging modalities on human visual perceptibility with the observed SSIM varies. The radiology images most clearly showed adversarial disturbances, with = 0.02 producing an already apparent, but very modest perturbation. At the same perturbation level for the ophthalmology and pathology photos, perturbations were nearly undetectable and became apparent with larger epsilon values. The authors examine black-box adversarial attacks on deep learning in healthcare imaging. [59]. They investigate three medical imaging sectors where deep learning algorithms are vulnerable. In all three datasets, perturbations calculated by FGSM showed lower SSIM than those calculated using PGD. This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We opted to document attacks in our subsequent analyses employing 0.02, as this represented the maximum perturbation level that remained effective across all applications and attack strategies. Additionally, it exhibited greater transferability than an epsilon value of 0.01 in the majority of the examined applications.
@@
Figure 9: In a range of black-box settings, such as target and surrogate pre-trained on ImageNet, target and surrogate at random initialized, target and replacement randomly initialized plus surrogate developed using a different and reduced dataset (d2/2), FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise. The aforementioned shows the average receiver operating characteristic curve (AUC) for both clean and black-box scenarios for each configuration. Diabetic retinopathy (DR) is misclassified in a red frame and correctly classified as referable or non-referable in a green frame. The adversarial noise  represents the difference between the initial and adversarial image.citeadversarial.bortsova2021.
@@
Figure 9 shows examples taken from the ophthalmology collection that demonstrate the transferability of assaults when the target and surrogate are randomly started and when both are pre-trained on ImageNet. Whether the target and surrogate models had the same architecture or a different one did not affect the findings.
5 Conclusion
@@
Diagnostic imaging AI systems based on computer vision are increasingly being used as models for classifying and segmenting diseases. The scaled-up use of medical imaging AI systems has given rise to serious safety issues because to DNNs’ susceptibility to adversarial samples. Recently, a number of ways have been put out to increase the efficacy of medical image defense tactics. Although numerous protection strategies have been put forth, there are still reservations over the use of medical deep learning techniques. This outcome arises from certain constraints within medical imaging, including a scarcity of high-quality image datasets and labeled data in comparison to more abundant datasets available for natural images. The efficacy of adversarial defenses may not be universally applicable in this context. The findings underscore the importance of exercising caution in the design of Deep Neural Networks (DNNs) for medical imaging and their subsequent real-world applications.
Declarations
Funding
A portion of the funding for this research comes from DOEd Grant P116Z220008 (1) and NSF Grant No. 2306109. The author(s) expresses all opinions, findings, and conclusions; the sponsor(s) does not necessarily agree with them.
Conflict of interest
The authors do not declare any conflicts of interest.
Ethics approval
There is no original research involving human subjects, animal subjects, or sensitive data in this review publication.
Consent to participate
The studies and data included in this review were sourced from publicly available and previously published literature.
Consent for publication
All authors listed on this review paper have reviewed and approved the final manuscript for submission to Machine Learning. Each author has contributed significantly to the research, writing, and revision of the paper. All authors have read and agreed to the content presented in this review, and their consent for publication is hereby provided.
Availability of data and materials
The data and materials used can be found in the references of this work.
Code availability
The review is based on a comprehensive analysis of publicly available literature, and any references to specific methodologies or algorithms are attributed to the respective original sources. We have strived to provide clear citations and references to enable readers to locate the original works for further examination.
Authors’ contributions
Conceptualization: A.B., M.A.A.N.; Methodology: A.B., M.A.A.N.; Formal analysis and investigation: A.B., M.A.A.N.; Writing - original draft preparation: A.B., M.A.A.N., R.G., A.R.; Writing - review and editing: M.A.A.N., A.R.; Funding acquisition: K.D.G; Resources: R.G.; Supervision: K.D.G., R.G. All authors have read and agreed to the published version of the manuscript.

### v2

@@ 3 Materials and Methods
This study examines various strategies to strengthen the usability of medical image analysis models against adversarial attacks. These strategies include techniques such as input denoising, adversarial training, adversary detection, image-level pre-processing, feature enhancement, and knowledge distillation. Each method tries to improve the aspects of models, which is necessary to the potential risks these attacks pose to clinical applications. Additionally, the research presents a new classification system for adversarial defense in medical imaging, categorized by application type, covering white-box, semi-white-box, and black-box attacks. By addressing the vulnerabilities in computerized diagnostic systems, this work contributes to the development of reliable and accurate healthcare technologies, ensuring consistent performance even under adversarial conditions.
@@ 3.1 Medical Adversarial Defensive Strategies
Defense models such as input denoising, input gradients regularization, and adversarial training have been created. The most recent attacks can typically completely or partially escape these countermeasures. In this case, a thorough investigation of adversarial attack and defense techniques in the field of medical image analysis is presented, including a family of techniques with a novel taxonomy based on the application situation [38]. A unified theoretical framework was built by the author for various adversarial attack and defense strategies in medical imaging applications. The white-box scenario is the main focus of most efforts on adversarial attack for medical image processing. These studies, in particular, with complete understanding of medical DNNs, focus on the adversarial vulnerability of computer-aided diagnosis models in diverse medical imaging applications. When executing adversary creation, the attacker might use the target diagnosis DNN as a locally deployed model.Additional academics have looked into the weaknesses of additional medical imaging tasks in addition to white-box adversarial attacks for medical classification tasks. These publications mostly focus on medical segmentation [31] [28]. For natural photos, semi-white-box (Gray-box) attacks have been extensively researched [39]. However, there are only a few academic publications [40], [41] that address this attack scenario for medical image processing. The semi-white-box adversarial approach typically consists of two stages: 1) To create adversarial instances against the target DNN model, the attacker trains a generative model. The attacker has complete access to the target model during training, including backward propagation gradients. 2) Instead of needing to know anything about the target model, as would be the case in a completely black-box scenario, the adversary generator can directly obtain adversarial examples against the target model with the input of authentic photos during the application stage. Currently used white-box adversarial attacks mostly call for numerous target model backward gradients. In other words, the attacker generates equivalent adversarial samples by treating the target DNN as the locally deployed model. However, because it requires a comprehensive understanding of the DNN model to attack, the white-box option can be unreliable in a real-world scenario. Comparatively, the general black-box scenario may be a better environment for simulating real-world adversarial attacks. There have been numerous proposals to investigate black-box assaults for natural images. The solutions outlined above either require numerous queries to the black-box model or depend on having complete understanding of the desired diagnosis model. However, in the majority of real-world scenarios, the attacker might not be able to directly access the target diagnosis model. The restricted black-box (no-box) configuration in particular might effectively represent the toughest (worst) situation for the real-world adversarial assault, even if it doesn’t call for querying the target blackbox DNN. The transferability [42] of adversarial cases across different DNN models is the fundamental determinant of the no-box attack. For instance, a no-box attacker can create adversarial images based on a surrogate model that has been deployed locally and that can be transferred straight to the target medical diagnosis systems. Restricted black-box adversarial attacks are more covertly dangerous for natural vision tasks, according to a well-known study. Given the significant dangers to the healthcare sector, a number of defense strategies have been put out to fend off medical hostile attacks. Numerous approaches have been put out to provide reliable deep learning-based systems for natural photos in response to the catastrophic failures brought on by adversarial instances [42]. Developing accurate computer-aided diagnosis models for clinical applications also helps millions of people receive trustworthy healthcare services. Investigating adversarially robust models in the context of medical image analysis is therefore important.
In [38], authors summarize the works on adversarial defense in the context of medical imagine analysis. In the context of medical pictures, various attack methods also serve as robustness evaluation criteria for adversarial defense, in addition to a considerable number of adversarial attack methods stated here.
@@ 3.3 Adversarial Detection
Adversary detection seeks to identify adversary cases from input examples during the application stage, as opposed to developing robustness during the training stage of computer-aided diagnosis models. In the field of biological image analysis, a variety of adversarial detection techniques have been put forth to prevent further misdiagnosis brought on by adversarial examples [36]. In particular, it is possible to think about medical adversarial detection as an anomaly detection issue that can be resolved by combining explainability approaches [46].
@@ 3.4 Image-level Pre-processing
3.5 Enhancement of Features
The discrepancy between the robustness of human and machine vision has been shown to be caused by adversarial examples being associated with non-robust features (extracted from specific patterns in the data distribution) [47]. Therefore, improving the feature representation is absolutely required for robust inference systems. In this study, we define the feature augmentation as the alteration of architectures or mapping functions. The robustness of medical classification models has been increased by the development of numerous feature improvement techniques [48].
@@ 3.6 Distillation of Knowledge
Knowledge distillation is a useful method for moving learnt information from a complicated (teacher) model to a simple (student) model in the machine learning field. As a result, self-distillation refers particularly to the scenario in which the network architectures for the teacher and student models are the same. Additionally, adversarial knowledge distillation, which transfers the adversarial resilience from the heavy teacher model to a light student model, has been extensively investigated for the natural imaging domain [49].
@@ 4 Result and Discussion
Four publicly available benchmark datasets are used in this study [38] to explore adversarial attack and defense for medical image processing (shown in Figure 4). 2) Messidor1 dataset of 1,200 eye fundus color numerical pictures for detecting diabetic retinopathy of four classes according to retinopathy grade; International Skin Imaging Collaboration (ISIC) dataset of 2,750 dermoscopic images of three different categories for skin lesion classification and segmentation. 3) The ChestX-ray 14 dataset consists of 112,120 frontalview X-ray images from 14 thorax diseases. 4) The COVID-19 database, which includes 21,165 chest X-ray images with segmentation-capable lung masks.
@@
Experiments are used by authors [50] to show: 1): Without affecting the classification accuracy of clean images, the SSAT module can considerably improve the adversarial robustness of the model. 2) The bulk of successful adversarial examples may be found and excluded by the UAD module. 3) When compared to other existing AI systems, their medical imaging AI solution (UAD + SSAT) minimizes adversarial risk. A publicly available dataset of retinal OCT images is used for the research. A publicly available dataset of retinal OCT images is used for the research. The most difficult threat used to assess class prediction performance is the ”white-box” setting.In all white-box attack conditions, the authors show that SSAT significantly outperforms alternative baselines while maintaining an equivalent or superior performance for clean picture classification. The authors show that the lowest adversarial risk for the new measure presented results from UAD complementing with SSAT. Regardless of the training techniques employed, it is evident that UAD-based systems consistently have reduced risks compared to those who do not.
@@
An AI-CAD model, in Figure 5, that intended to alter the diagnosis-sensitive contents of images (by adding or deleting malignant tissue) was first trained, and then it was evaluated on the adversarial images produced by the GAN model. The reader study looked at how well human specialists could visually identify the images created by the GAN. Authors [51] of a research paper conducted a study to assess an AI-CAD model’s responses to adversarial attacks on GAN-generated mammography pictures by introducing malignant tissue into healthy photos and removing cancerous tissue from cancer-affected images. Additionally, they assessed the effectiveness of experienced radiologists in recognizing these types of antagonistic pictures visually both without and after an instructional intervention. The University of Pittsburgh Medical Center provided the authors with 1284 women for their study cohort, and they also collected 4346 mammography pictures for this cohort. In this cohort, 366 patients had breast cancer that was biopsy-proven to be malignant while 918 patients had breast cancer that was evaluated as benign (including benign signs). Following training, researchers evaluated the model’s classification accuracy in comparison to both the original genuine test data and its equivalent GAN-generated adversarial counterparts. To generate the adversarial images, we created two GAN-trained U-Net23 models. The GAN generators, which were trained to create fake positive mammography images from the falsely negative mammogram photos and fake negative mammogram images from the falsely positive mammogram images, respectively, updated the mammogram images in the test set with flipped labels. The two distinct resolutions of this image indicate the categorization impacts of the AI-CAD model on the test data. The authors begin by outlining the findings from the high-resolution (1728 1408) photographs. As can be observed, the AI-CAD model obtained an AUC of 0.82 on the test set, which consisted of 364 genuine negative samples and 74 real positive samples. The test set’s equivalent adversarial GAN-generated images (with flipped labels) yielded an AUC of 0.94. These two AUC values show that the adversarial collection of photos generally succeeded in deceiving the AI-CAD model. Additionally, when calculating classification accuracy using a threshold of 0.5, 44 out of the 74 real positive images (or 59.5%) were correctly labeled as positive cases; however, only 42 out of the 44 fake negative images (note that their labels were negative due to the flipping) were categorized as negative cases, indicating that 95.5% (42 out of 44 cases) of the GAN-generated adversarial samples successfully deceived the system.
@@
Figure 6: Between the original images in the test sets and the adversarial examples produced using FGSM or PGD with varied perturbation degrees, the Mean Structural Similarity Index Measure (SSIM) was calculated. The SSIM values shown by the points are averaged over two model architectures (Inception-v3 and Densenet-121).[52]
@@
Figure 6 displays the mean SSIM values for FGSM and PGD assaults across all photos. The Supplementary Material includes the SSIM data for every model. As can be shown, the effects of the identical disturbance applied to various imaging modalities on human visual perceptibility with the observed SSIM varies. The radiology images most clearly showed adversarial disturbances, with = 0.02 producing an already apparent, but very modest perturbation. At the same perturbation level for the ophthalmology and pathology photos, perturbations were nearly undetectable and became apparent with larger epsilon values. Black-box adversarial assaults on deep learning in medical imaging are studied by the authors [52]. They investigate three medical imaging sectors where deep learning algorithms are vulnerable. In all three datasets, perturbations calculated by FGSM showed lower SSIM than those calculated using PGD. This is a predicted outcome given that PGD optimizes perturbations based on their size and influence on model predictions.We decided to report attacks using for our further investigations utilizing = 0.02, since this was the maximum degree of disturbance that was still for all applications and assault strategies, and it had more transferability than an epsilon of 0.01 in most cases of the applications examined.
@@
Figure 7: FGSM (=0.02) was used to create the original images, adversarial images, and corresponding adversarial noise in a variety of black-box settings, including target and surrogate pre-trained on ImageNet, target and surrogate randomly initialized, target and surrogate randomly initialized plus surrogate developed using a different and reduced dataset (d2/2). The average receiver operating characteristic curve (AUC) for each configuration for both clean and black-box scenarios is shown above. Diabetic retinopathy (DR) is correctly classified as referable or non-referable in a green frame and incorrectly classified in a red frame. The difference between the original and the adversarial image is represented by the adversarial noise [52].
@@
Examples from the ophthalmology dataset are shown in Figure 7 to demonstrate attack transferability when the target and surrogate are both pre-trained on ImageNet and when both are initialized randomly. The instances where the target and surrogate models had the same or different architectures had no effect on any of the aforementioned results.
5 Conclusion and Discussion
@@
The study identifies the impact of adversarial defense strategies in medical image analysis using datasets like Messidor1, ISIC, ChestX-ray14, and COVID-19. The SSAT module significantly improves model robustness against adversarial attacks while maintaining accuracy on clean images. Combined with the UAD module, it efficiently detects and removes adversarial examples, minimizing risks in upcoming applications. Diagnostic imaging AI systems based on computer vision are increasingly being used as models for classifying and segmenting diseases. The scaled-up use of medical imaging AI systems has given rise to serious safety issues because to DNNs’ susceptibility to adversarial samples. Recently, a number of ways have been put out to increase the efficacy of medical image defense tactics. Although numerous protection strategies have been put forth, there are still reservations over the use of medical deep learning techniques. This is a result of some limitations in medical imaging, such as a dearth of high-quality picture datasets and labeled data as compared to other high-quality natural image datasets. Adversarial defenses might not have a limited impact. The results of the study highlight the need for greater care in designing DNNs for medical imaging and their practical applications.
Figure 8: A demonstration of how adversarial attacks against different medical AI systems could be carried out without requiring any openly dishonest data manipulation [53].